"James invested some money at 10% interest. James also invested $90 more than 3 times that amount at 8%. How much is invested at each rate if James receives $798.38 in interest after one year?"
Found 2 solutions by tehreem, MathTherapy:
Answer by tehreem(2)


let the investment be $x
$x at 10% interest is x(1+0.1) = 1.1x
$(90+3x) at 8% interest = 1.08 (90 +3x) = 97.2 + 3.24x
if the total amount he receives by the end of the year is 798.38 , then:
1.1x + 97.2 + 3.24x = 798.38
4.34x = 798.38 - 97.2
4.34x = 701.18
x = 162

(now substitute the value of x )
1.1 x 162= $178 at 10 % interest rate
97.2 + 3.24(162) = $622 at 8% interest rate


Answer by MathTherapy(10557)


"James invested some money at 10% interest. James also invested $90 more than 3 times that amount at 8%. How much is invested at each rate if James receives $798.38 in interest after one year?"